MEAN VALUES OF THE LOGARITHMIC DERIVATIVE OF

THE ZETA FUNCTION AND THE GUE HYPOTHESIS

David W. Farmer

Abstract. The GUE Hypothesis, which concerns the distribution of zeros of the Riemann zeta-function, is used to evaluate some integrals involving the logarithmic derivative of the zeta-function. Some connections are shown between the GUE Hypothesis and other conjectures.

Typeset by AMS-TEX